OpenStudy (anonymous):

I have 2 credit cards the total available on both is $2,300. The interest on one is 9.8% and the other 12.4% after year. what is the dollar amount for each card? Someone helped me with it but just gave the answer I need to see steps so I can understand it better any help would be great Thanks.

OpenStudy (anonymous):

your question doesn't make sense to me

hero (hero):

I was going to say something about that, but I wasn't sure. Does EACH card have 2,300 or is the total combined balance 2300? I think the question "What is the dollar amount for each card" needs to be expressed more precisely. Something is wrong with it.

OpenStudy (anonymous):

I have 2 credit cards the total available on both is $2,300. The interest on one is 9.8% and the other 12.4% after year. what is the dollar amount for each card? thats all it says ill just pass on it thanks guys here is what I have a+c=2,300 9.8(a)+12.4(c)=2300 98(a)+124(c)=2300

OpenStudy (mathmate):

You may have skipped one line of the question. Does it say somewhere that the interest on both cards at the end of one year is the same?